Write each of the following as decimals:
12 / 5 ????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

12 / 5 = (10 + 2) / 5
= 10 / 5 + 2 / 5
= 2 + 0.4
= 2.4

Thanks